A small object begins a free-fall from a height of H 865 m at to 0 s. After τ-2.60 s, a second small object is launched vertically up from the ground with an initial velocity of vo-41.2 m/s. At what height from the ground will the two objects first meet? height:
Answer #1

ibistance travelled by the freely falling object - using s=uto + 1 gt2 S = 0 + LX98 x(2.60)? S = 33.124 m Distance cover will